OpenStudy (anonymous):

Rationalize the denominator of the expression. Then simplify your answer. 2/(7 – √3). i got an answer of (14+2√3)/46 but its wrong

OpenStudy (anonymous):

they want you to cancel a two top and bottom i bet

OpenStudy (anonymous):

\[\frac{7+\sqrt{3}}{23}\]

OpenStudy (anonymous):

hmm, i'll try that

OpenStudy (anonymous):

yea they did, how do you know when to do it? i've come across problems where they don't want you to for some reason.

OpenStudy (anonymous):

(9-6√3)/12. can i apply that to this as well?

OpenStudy (anonymous):

it was a guess when you said it didn't like the answer.. generally you should write fraction in reduced form. so even if you had \[\frac{14+2x}{46}\] you would probably write \[\frac{7+x}{23}\]

OpenStudy (anonymous):

yes it would e \[\frac{3-2\sqrt{3}}{4}\]

OpenStudy (anonymous):

You got it right...you just didn't do the simplification part \[(14 + 2\sqrt{3})/4 = (7 + \sqrt{3})/23\]

OpenStudy (phi):

you asked "yea they did, how do you know when to do it? " The question states "Then simplify your answer." Simplify, in this case, means factor out any common multiples In your answer, each term can be divided by 2 (in both the numerator and denominator).